# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 3.1 Bounding the topology of the bordism
irrespective of the choice of bordism realization $\mathcal{B}_{n+1}$. In general we will expect $\mathcal{B}_{n+1}$ to be connected and not enclosing any volume, so that $b_{0}(\mathcal{B}_{n+1})=1$ and $b_{n+1}(\mathcal{B}_{n+1})=0$, which further simplifies computations. Furthermore, for most of the examples considered, the torsion parts will be trivial.
Denoting $\vec{b}(\mathcal{X}_{n})=(b_{0}(\mathcal{X}_{n}),\dots,b_{n}(\mathcal{X}_{n}))$, we can illustrate the above relations (3.8) and (3.9) with some examples:
@@
  • $\mathcal{C}_{3}=\mathbb{T}^{3}$: $\vec{b}(\mathbb{T}^{3})=(1,3,3,1)$, so that $\vec{b}(\mathcal{B}_{4})=(1,\beta_{1},\beta_{2},\beta_{3},0)$, with $\beta_{1},\,\beta_{2}\,,\beta_{3}\geq 0$ and $\beta_{1}+\beta_{2}\geq 3$. Note that (though in this case just 2) the number of possibilities saturating the inequalities is not unique. The first one of them, $\vec{b}(\mathcal{B}_{4})=(1,2,1,0,0)$, corresponds with the straightforward bordism $\mathcal{B}_{4}=\mathbb{T}^{2}\times\mathbb{D}^{2}$, while $\vec{b}(\mathcal{B}_{4})=(1,1,2,0,0)$ does not immediately correspond to any known 4-manifold with boundary. We will come back to this example is section 3.3 with a more involved bordism.
@@
[…]
@@ 3.2 Morse-Bott inequalities for manifolds with boundary
Once topological information for the bordism $\mathcal{B}_{n+1}$ has been obtained, it is time to relate it with the number and type of necessary topology changes $\mathcal{C}_{n}\to\cdots\to\emptyset$. An important tool to do this will be the ansatze (2.8) and (2.9) for the BoN solutions, such that we can split the bordism and angular part of our solutions, and the radial direction $\rho:\mathcal{B}_{n+1}\to(-\infty,0]$, [^7] For our purposes in this section, we will want $\rho$ to grown towards the tip of the bordism, so we will take this definition rather than $\rho:\mathcal{B}_{n+1}\to[0,+\infty)$. which indicates the distance to the BoN/EotW wall. Note the following two properties of $\rho$:
  • As given in (2.9), for $\rho\to\infty$, we can split $h_{\alpha\beta}^{\mathcal{B}}(y)\mathrm{d}y^{\alpha}\mathrm{d}y^{\beta}\approx\mathrm{d}\rho^{2}+{h}^{\mathcal{C}}_{\bar{\alpha}\bar{\beta}}\mathrm{d}y^{\bar{\alpha}}\mathrm{d}y^{\bar{\beta}}$, so that in this limit $\mathcal{B}_{n+1}$ locally looks like ×$\mathbb{R}\times\mathcal{C}_{n}$. This means that in this regime $\mathrm{d}\rho\neq 0$, as $h^{\mathcal{B}}_{\alpha\beta}$ is non-degenerate.
@@
The above properties mean that $\rho$ is a Morse-Bott function [^9] In more precise words, given $\mathcal{B}_{n+1}$ a $(n+1)$-dimensional compact manifold with boundary $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, a smooth function $\rho:\mathcal{B}_{n+1}\to\mathbb{R}$ is called Morse-Bott if the set of critical points ${\rm Crit}(f)=\{p\in\mathcal{B}_{n+1}:\,\mathrm{d}f(p)=0\}$ is a disjoint union of connected sub-manifolds of $\mathcal{B}_{n+1}-\mathcal{C}_{n}$, with each connected component being non-degenerate, and $\left.\rho\right|_{\mathcal{C}_{n}}:X_{n}\to\mathbb{R}$ is also Morse-Bott. Note that in principle this would also allow $\mathcal{C}_{n}$ to be a manifold with boundary., which will allow us to use the so-called Morse-Bott inequalities (for manifolds with boundary) ORITA_2018 . In what follows we introduce the needed mathematical machinery in order to use them.
Take for this $\rho$ to be Morse-Bott function. Then a critical point of the boundary $p\in{\rm Crit}(\rho|_{\mathcal{C}_{n}})$ is said to be type $N$ if $\langle\mathrm{d}\rho|_{\mathcal{C}_{n}}(p),n(p)\rangle<0$, where $n(p)\in T_{p}\mathcal{B}_{n+1}$ is an outward [^10] In the sense of pointing out of the boundary $\mathcal{C}_{n}$ in a direction opposite of the interior of $\mathcal{B}_{n+1}$, rather than an inward normal vector pointing towards the interior of $\mathcal{B}_{n+1}$. These are respectively represented by $n$ and $m$ in Figure 4. normal vector to the boundary at $p$. We then define the following submanifolds and connected components:
@@
We will denote by $c_{l}$ and $d_{s}$ the respective dimensions of $\mathfrak{C}_{j}$ and $\Gamma_{s}$, and by $\lambda_{j}$ and $\mu_{s}$ the index of $\rho$ and $\rho|_{\mathcal{C}_{n}}$ over them (i.e. the number of negative eigenvalues of its Hessian at each point of the critical submanifold). These different submanifolds and concepts are depicted in Figure 4.
Figure 4: Sketch of the bordism $\mathcal{B}_{3}$ for the genus-2 Riemann surface $\Sigma_{2}$. The different critical submanifolds of the Morse-Bott function $\rho$ are depicted in black, $\mathfrak{C}_{1}\simeq\{p\}$ ($c_{1}=0$) and $\mathfrak{C}_{2}\simeq\mathbb{S}^{1}$ ($c_{2}=1$), both with indices $\lambda_{1}=\lambda_{2}=0$. Also depicted are two outward and inward normal vectors to the boundary, $n$ and $m$, as well as the vector field $\mathrm{d}\rho^{\flat}=h^{\mathcal{B}\,\alpha\beta}\partial_{\alpha}\rho\partial_{\beta}=\partial^{\alpha}\rho\partial_{\alpha}$ associated to $\rho$, in blue. Depicted in green are the internal 2-manifold at two fixed values of $\rho$.
@@
for some $\mathsf{R}(t)\in\mathbb{Z}_{\geq 0}[t]$. Note that evaluating at $t=-1$, we have $\chi(\mathcal{B}_{n+1})=\chi(\mathcal{C}_{n})+\sum_{j=1}^{l}(-1)^{\lambda_{j}}\chi(\mathfrak{C}_{j})$, with $\chi(\mathcal{M})$ the Euler characteristic of $\mathcal{M}$.
@@
Topology changes occur when internal cycles shrink over critical submanifolds $\mathfrak{C}$ of $\rho$, with $\mathrm{d}\rho=0$ indicating that such cycle no longer “propagates” as we move towards the BoN/EotW brane. The dimensionality of the “disappearing” cycle is given by $c+\lambda-1$, where $c=\dim\mathfrak{C}$ and $\lambda$ its critical index, as $\lambda-1$ dimensions shrink (one has to discount the radial direction) over the $c$-dimensional base the cycle is fibered over. Furthermore, apart from this $(c+\lambda-1)$-cycle $\mathcal{A}$ that shrinks (what we could interpret as a cobordism for $\mathcal{A}$ into nothing), the dual $(n-c-\lambda+1)$-cycle $\mathcal{A}^{\prime}$ also disappears, not shrinking per se but being interpreted now as a chain with boundary at $\mathcal{C}$. This results in a reduction of the homology groups $H_{c+\lambda-1}(\rho^{-1}(r);\mathbb{Z})$ and $H_{n-c-\lambda+1}(\rho^{-1}(r);\mathbb{Z})$ for $r$ smaller and larger than $r_{*}$, with $r_{*}$ the location of the critical point, which translates into a lowering (not necessarily by a unit) of the respective Betti numbers and/or torsion ranks. We will revisit this cycle shrinking/transforming into a chain with boundary interpretation when discussing flux potentials and their fate after topology changes in section 4.2.
@@
Analogously, it could be the case that the topology change results in the creation of some cycle (and its dual), as occurs in Figure 3(b) or by the nucleation of a baby Universe that might reattach later or not, but both of these cases the cobordism into nothing will feature more topology changes than necessary.
@@
The number $l$ of critical submanifolds will corresponds to the number of topology changes (including the ultimate collapse into nothing), so impossibility to satisfy (3.13) with $l=1$ means an intermediate topology change must occur. Note that the index of the critical submanifold is given by $\lambda+c-1=n$, as all dimensions must close at the top of the BoN/EotW brane.
@@
Inequalities 3.13 only consider the free part of the homology groups (i.e. the Betti numbers), with the torsion ranks not entering into account. While most of the examples considered will not feature torsion, one would expect that bordisms with torsion will probably require additional topology changes in order to get rid of this topological charge. Unfortunately, to our knowledge there are not results concerning Morse-Bott inequalities for manifolds with boundary and torsion in the Mathematics literature. The closest result are the Morse-Pitcher inequalities Pitcher for closed manifolds without boundary (but with torsion). Having $\rho:\mathcal{M}_{n}\to\mathbb{R}$ a Morse function (this is, critical loci will be isolated points rather than general submanifolds, as in Morse-Bott functions), then
@@ 3.2.1 Some easy examples
so that one can take $\mathsf{R}(t)=t(1+t)^{n-1}$, which means that on principle intermediate topology changes are not necessary for BoN with compact manifold $\mathbb{T}^{n}$. We will see in section 4.3 more involved bordisms of toroidal compactifications which include additional topology changes.
@@
such that the only critical submanifold of $\hat{\rho}$ is ^=×$\hat{\mathfrak{C}}=\mathfrak{C}\times Z_{n-m}$. Finally, using $\mathsf{P}_{t}(A\times B)=\mathsf{P}_{t}(A)\mathsf{P}_{t}(B)$, we obtain
@@
As the Poincaré polynomial always has non-negative coefficients, $\mathsf{R}(t)\mathsf{P}_{t}(\mathcal{Y}_{n-m})\in\mathbb{Z}_{\geq 0}[t]$, so that on principle there is no need for intermediate topology changes in $\mathcal{C}_{n}=\mathcal{X}_{M}\times\mathcal{Y}_{n-m}$ going to nothing. As we will see in section 4.4, this does not extend to non-trivial fibrations.
@@ 3.3 Non-minimal bordisms and defects
In the above sections only the global topology (actually only the homology groups) of the bordism $\mathcal{B}_{n+1}$ was considered, assuming from the start that the relevant bordism groups $\Omega^{\mathsf{g}}_{n}$ vanished. However, as explained in section 1, in order for this to happen there might be extra ingredients or defects that must be added to our theory, such that $\Omega_{n}^{\mathsf{g}}\neq 0\to\Omega^{\mathsf{g}+{\rm def.}}_{n}=0$. In a general way, $\Omega_{k}^{\mathsf{g}}\neq 0$ implies the existence of a $D-k-1$-global symmetry in our theory, with associated conserved topological charge/invariant McNamara:2019rup
@@
where $G_{k}$ is some associated closed $k$-form. Note that by definition $Q$ is independent of the bordism representative. Now, in order to kill this bordism invariant, a $D-k-1$ defect is needed. For the case of oriented bordisms in the presence of $A_{p}$ forms, $\Omega_{p+1}^{SO,\,A_{p}}\simeq\mathbb{Z}$, represented by the magnetic charge $Q$ above. The $D-(p+1)-1=D-p-2$ defects are given by the D$(D-p-3)$-branes on whose world-volume the magnetic dual gauge potential $\tilde{A}_{D-p-2}$ lives (see McNamara:2019rup for more details on this).
@@
Figure 5: Illustrations of thin and thick walls in different bordisms, respectively for $\mathcal{C}_{2}=\mathbb{T}^{2}$ (subfigure 5(a)) and $\mathcal{C}_{1}=\mathbb{S}^{1}$ (subfigure 5(b)). In the thin wall case the $(d+n-k-1=d)$-defect simply extends over $d-1$ spacetime directions and wrapping along the $(n-k=2-1=1)$-cycle $\tilde{a}_{1}$ dual to the 1-cycle $a_{1}$ that is shrinking. In the thick wall case the analogous $d$-defect extends over $d-1$ spacetime directions and wraps a 1-cycle $A_{1}$ of $\mathcal{B}_{2}\simeq\mathbb{T}^{2}-\mathbb{D}^{2}$. This second example can be found in section 7.3 from Debray:2023yrs .
@@
In general, when compactifying on $\mathcal{C}_{n}$ to $d=D-n$ dimensions, we will also have internal $k$-cycles over which topological charges can be defined. The $d+n-k-1$-defects associated with $\{\Omega_{k}^{\mathsf{g}}\}_{k=0}^{n}$ will allow the associated $k$-cycles to go to nothing in the intermediate topology changes described in section 3.2. As at the different sides of the topology change (3.1) the compactification manifold $\mathcal{C}_{n}^{(k)}\to\mathcal{C}_{n}^{(k+1)}$ will be different, the lower $d$-dimensional EFT will change, with the $d+n-k-1$-defects being seen from the macroscopic point of view as codimension-1 $d-1$-domain walls, which are localized at the topology change loci. Depending on how the additional $n-k$ directions are wrapped, we can have the following two types of domain walls:
@@
  • Thin domain walls: The $d+n-k-1$-defect extends over $d-1$ of the macroscopic directions, while wrapping the $n-k$-cycle dual to the $k$-cycle in $\mathcal{C}_{n}$ supporting the topological charge, and that the defect allows to shrink to nothing. Note that after the topology change the dual $n-k$-cycle no longer exists either, so the domain wall also acts as a boundary of both cycles. An illustration of this is depicted in Figure 5(a), and will found in explicit examples in section 4.2.
@@
  • Thick domain walls: The $d+n-k-1$-defect extends over $d-1$ macroscopic directions (perpendicular to the bordism radial direction $\rho$), while wrapping a $n-k$-cycle in the bordism $\mathcal{B}_{n+1}$, in such a way that the cycle also extends in the $\rho$ direction (for a finite interval), with the $n-k$-cycle not corresponding to any specific cycle in any of the $\mathcal{C}_{n}^{(k)}$ slices. Note that in this case the localization occurs around said cycle, which by definition is non-contractible, unlike those of $\mathcal{C}_{n}^{(k)}$ once embedded in $\mathcal{B}_{n+1}$.
@@
    This is illustrated in Figure 5(b), with an example for this disposition being found in section 7.3 of Debray:2023yrs (albeit not in the context of an explicit BoN construction), the non-trivial 1-cycle $A_{1}\subset\mathcal{B}_{2}$ has a monodromy that kills the $\mathbb{Z}_{3}$ factor of $\Omega_{1}^{{\rm Spin-Mp}(2,\mathbb{Z})}\simeq\mathbb{Z}_{8}\oplus\mathbb{Z}_{3}$ by performing a reflection $g\to g^{-1}$ on the generator of $\mathbb{Z}_{3}=\langle g\rangle$ (we refer to the original, Debray:2023yrs for more details). This is equivalent to having a 8-defect spanning between the two critical points (in black in Figure 5(b)) associated to $A_{1}$, resulting in a thick (7+1)-domain wall. [^11] We thank Markus Dierigl and Miguel Montero for referencing a specific example of this class to us.
@@
The defects considered here in general are of stringy origin and will backreact on the spacetime geometry, resulting in a loss of the smooth geometry (e.g. conical defects and angular deficits). We expect that, in some of these cases, the topological description of our bordism to hold in a way that can be smoothed-out and the Morse-Bott inequalities hold.

### v2

@@ 3.1 Bounding the topology of the bordism
irrespective of the choice of bordism realization $\mathcal{B}_{n+1}$. In general we will expect $\mathcal{B}_{n+1}$ to be connected and not enclosing any volume, so that $b_{0}(\mathcal{B}_{n+1})=1$ and $b_{n+1}(\mathcal{B}_{n+1})=0$, which further simplifies computations.
Unlike the free part, the torsion part of the homology is not determined (up to isomorphisms) by the torsion rank. However, using the Fundamental Theorem of finite abelian groups one can show that $T_{k}(\mathcal{C}_{n})\simeq S_{k}\oplus S^{\prime}_{n-k-1}$, with $S_{k}\leq T_{k}(\mathcal{B}_{n+1})$ and $S^{\prime}_{n-k-1}\leq T_{n-k-1}(\mathcal{B}_{n+1})$, though we will not make use of this relation along the paper. Furthermore, for most of the examples considered, the torsion parts will be trivial.
Although along this paper we will focus on bordisms to nothing, such that $\mathcal{C}_{n}$ has unique component, the above relations hold in generals, so in principle one could apply them for $\partial\mathcal{B}_{n+1}=\mathcal{C}_{n}\sqcup\overline{\mathcal{C}}_{n}^{\prime}$, with bordisms between two compactifications.
Now, denoting $\vec{b}(\mathcal{X}_{n})=(b_{0}(\mathcal{X}_{n}),\dots,b_{n}(\mathcal{X}_{n}))$, we can illustrate the above relations (3.8) and (3.9) with some examples:
@@
  • $\mathcal{C}_{3}=\mathbb{T}^{3}$: $\vec{b}(\mathbb{T}^{3})=(1,3,3,1)$, so that $\vec{b}(\mathcal{B}_{4})=(1,\beta_{1},\beta_{2},\beta_{3},0)$, with $\beta_{1},\,\beta_{2}\,,\beta_{3}\geq 0$ and $\beta_{1}+\beta_{2}\geq 3$. Note that (though in this case just 2) the number of possibilities saturating the inequalities is not unique. The first one of them, $\vec{b}(\mathcal{B}_{4})=(1,2,1,0,0)$, corresponds to the straightforward bordism $\mathcal{B}_{4}=\mathbb{T}^{2}\times\mathbb{D}^{2}$, while $\vec{b}(\mathcal{B}_{4})=(1,1,2,0,0)$ does not immediately correspond to any known 4-manifold with boundary. We will come back to this example is section 4.3 with a more involved bordism.
@@
  • We finally take the lens space $\mathcal{C}_{3}=L(p;q)$, with homology groups $\mathbb{Z}$, $\mathbb{Z}_{p}$, 0 and $\mathbb{Z}$, which unlike the previous examples has torsion. We find that $\vec{b}(\mathcal{B}_{4})=(1,\beta_{1},\beta_{2},\beta_{3},0)$ with $\beta_{i}\geq 0$ for i=1, 2, 3i=1,\,2,\,3, while $t_{1}(\mathcal{B}_{4})\geq 1$ (the rest of $\mathcal{B}_{4}$ torsion ranks can be null).
@@ 3.2 Morse-Bott inequalities for manifolds with boundary
Once topological information for the bordism $\mathcal{B}_{n+1}$ has been obtained, it is time to relate it with the number and type of necessary topology changes $\mathcal{C}_{n}\to\cdots\to\emptyset$. An important tool to do this will be the ansätze (2.8) and (2.9) for the BoN solutions, such that we can split the bordism and angular part of our solutions, and the radial direction $\rho:\mathcal{B}_{n+1}\to(-\infty,0]$, [^9] For our purposes in this section, we will want $\rho$ to grown towards the tip of the bordism, so we will take this definition rather than $\rho:\mathcal{B}_{n+1}\to[0,+\infty)$. which indicates the distance to the BoN/EotW wall. Note the following two properties of $\rho$:
  • As given in (2.9), for $\rho\to\infty$, we can split $h_{\alpha\beta}^{\mathcal{B}}(y)\mathrm{d}y^{\alpha}\mathrm{d}y^{\beta}\approx\mathrm{d}\rho^{2}+{h}^{\mathcal{C}}_{\bar{\alpha}\bar{\beta}}\mathrm{d}y^{\bar{\alpha}}\mathrm{d}y^{\bar{\beta}}$, so that in this limit $\mathcal{B}_{n+1}$ locally looks like $\mathbb{R}\times\mathcal{C}_{n}$. This means that in this regime $\mathrm{d}\rho\neq 0$, as $h^{\mathcal{B}}_{\alpha\beta}$ is non-degenerate.
@@
The above properties mean that $\rho$ is a Morse-Bott function [^11] In more precise words, given $\mathcal{B}_{n+1}$ a $(n+1)$-dimensional compact manifold with boundary $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, a smooth function $\rho:\mathcal{B}_{n+1}\to\mathbb{R}$ is called Morse-Bott if the set of critical points ${\rm Crit}(f)=\{p\in\mathcal{B}_{n+1}:\,\mathrm{d}f(p)=0\}$ is a disjoint union of connected sub-manifolds of $\mathcal{B}_{n+1}-\mathcal{C}_{n}$, with each connected component being non-degenerate, and $\left.\rho\right|_{\mathcal{C}_{n}}:X_{n}\to\mathbb{R}$ is also Morse-Bott. Note that in principle this would also allow $\mathcal{C}_{n}$ to be a manifold with boundary., which will allow us to use the so-called Morse-Bott inequalities (for manifolds with boundary) Orita (2018). In what follows we introduce the needed mathematical machinery in order to use them.
Take for this $\rho$ to be a Morse-Bott function. Then a critical point of the boundary $p\in{\rm Crit}(\rho|_{\mathcal{C}_{n}})$ is said to be type $N$ if $\langle\mathrm{d}\rho|_{\mathcal{C}_{n}}(p),n(p)\rangle<0$, where $n(p)\in T_{p}\mathcal{B}_{n+1}$ is an outward [^12] In the sense of pointing out of the boundary $\mathcal{C}_{n}$ in a direction opposite to the interior of $\mathcal{B}_{n+1}$, rather than an inward normal vector pointing towards the interior of $\mathcal{B}_{n+1}$. These are respectively represented by $n$ and $m$ in Figure 4. normal vector to the boundary at $p$. We then define the following submanifolds and connected components:
@@
We will denote by $c_{j}$ and $d_{s}$ the respective dimensions of $\mathfrak{C}_{j}$ and $\Gamma_{s}$, and by $\lambda_{j}$ and $\mu_{s}$ the Morse index of $\rho$ and $\rho|_{\mathcal{C}_{n}}$ over them (i.e. the number of negative eigenvalues of its Hessian at each point of the critical submanifold). These different submanifolds and concepts are depicted in Figure 4.
Figure 4: Sketch of the bordism $\mathcal{B}_{3}$ for the genus-2 Riemann surface $\Sigma_{2}$. The different critical submanifolds of the Morse-Bott function $\rho$ are depicted in black, $\mathfrak{C}_{1}\simeq\{p\}$ ($c_{1}=0$) and $\mathfrak{C}_{2}\simeq\mathbb{S}^{1}$ ($c_{2}=1$), both with indices $\lambda_{1}=\lambda_{2}=2$. Also depicted are two outward and inward normal vectors to the boundary, $n$ and $m$, as well as the vector field $\mathrm{d}\rho^{\flat}=h^{\mathcal{B}\,\alpha\beta}\partial_{\alpha}\rho\partial_{\beta}=\partial^{\alpha}\rho\partial_{\alpha}$ associated to $\rho$, in blue. Depicted in green are the internal 2-manifolds at two fixed values of $\rho$.
@@
for some $\mathsf{R}(t)\in\mathbb{Z}_{\geq 0}[t]$. Note that evaluating at t=1t=-1, we have $\chi(\mathcal{B}_{n+1})=\chi(\mathcal{C}_{n})+\sum_{j=1}^{l}(-1)^{\lambda_{j}}\chi(\mathfrak{C}_{j})$, with $\chi(\mathcal{M})$ the Euler characteristic of $\mathcal{M}$.
@@
Topology changes occur when internal cycles $\mathcal{A}\subseteq\mathcal{C}_{n}$ shrink over critical submanifolds $\mathfrak{C}$ of $\rho$, with $\mathrm{d}\rho=0$ indicating that such cycle no longer “propagates” as we move towards the BoN/EotW brane. The dimensionality of the “disappearing” cycle is given by $c+\lambda-1$, where $c=\dim\mathfrak{C}$ and $\lambda$ its critical index, as $\lambda-1$ dimensions shrink (one has to discount the radial direction) over the $c$-dimensional base the cycle is fibered over. Furthermore, apart from this $(c+\lambda-1)$-cycle $\mathcal{A}$ that shrinks (what we could interpret as a cobordism for $\mathcal{A}$ into nothing), the dual $(n-c-\lambda+1)$-cycle $\mathcal{A}^{\prime}$ also disappears, not shrinking per se but being interpreted now as a chain with boundary at $\mathcal{C}$. This results in a reduction of the homology groups $H_{c+\lambda-1}(\rho^{-1}(r);\mathbb{Z})$ and $H_{n-c-\lambda+1}(\rho^{-1}(r);\mathbb{Z})$ for $r$ smaller and larger than $r_{*}$, with $r_{*}$ the location of the critical point. This translates into a lowering (not necessarily by a unit) of the respective Betti numbers and/or torsion ranks. We will revisit this cycle shrinking/transforming into a chain with boundary interpretation when discussing flux potentials and their fate after topology changes in section 4.2.
@@
Analogously, it could be the case that the topology change results in the creation of some cycle (and its dual), as occurs in Figure 3(b), but for these cases the cobordism into nothing will feature more topology changes than necessary.
@@
The number $l$ of critical submanifolds will correspond to the number of topology changes (including the ultimate collapse into nothing), so impossibility to satisfy (3.13) with $l=1$ means an intermediate topology change must occur. Note that the index of the final critical submanifold is given by $\lambda+c-1=n$, as all dimensions must close at the top of the BoN/EotW brane.
@@
Inequalities 3.13 only consider the free part of the homology groups (i.e. the Betti numbers), with the torsion ranks not entering into account. While most of the examples considered will not feature torsion, one would expect that bordisms with torsion will probably require additional topology changes in order to get rid of this topological charge. Unfortunately, to our knowledge there are not results concerning Morse-Bott inequalities for manifolds with boundary and torsion in the Mathematics literature. The closest result are the Morse-Pitcher inequalities Pitcher (1958) for closed manifolds without boundary (but with torsion). Having $\rho:\mathcal{M}_{n}\to\mathbb{R}$ a Morse function (this is, critical loci will be isolated points rather than general submanifolds, as in Morse-Bott functions), then
@@ 3.2.1 Some easy examples
so that one can take $\mathsf{R}(t)=t(1+t)^{n-1}$, which means that in principle intermediate topology changes are not necessary for BoN with compact manifold $\mathbb{T}^{n}$. We will see in section 4.3 more involved bordisms of toroidal compactifications which include additional topology changes.
@@
such that the only critical submanifold of $\hat{\rho}$ is $\hat{\mathfrak{C}}=\mathfrak{C}\times\mathcal{Y}_{n-m}$. Finally, using $\mathsf{P}_{t}(A\times B)=\mathsf{P}_{t}(A)\mathsf{P}_{t}(B)$, we obtain
@@
As the Poincaré polynomial always has non-negative coefficients, $\mathsf{R}(t)\mathsf{P}_{t}(\mathcal{Y}_{n-m})\in\mathbb{Z}_{\geq 0}[t]$, so that in principle there is no need for intermediate topology changes in $\mathcal{C}_{n}=\mathcal{X}_{M}\times\mathcal{Y}_{n-m}$ going to nothing. As we will see in section 4.4, this does not extend to non-trivial fibrations.
@@ 3.3 Non-minimal bordisms and defects
In the above sections only the global topology (actually only the homology groups) of the bordism $\mathcal{B}_{n+1}$ was considered, assuming from the start that the relevant bordism groups $\Omega^{\mathsf{g}}_{n}$ vanished. However, as explained in section 1, in order for this to happen there might be extra ingredients or defects that must be added to our theory, such that $\Omega_{n}^{\mathsf{g}}\neq 0\to\Omega^{\mathsf{g}+{\rm def.}}_{n}=0$. In a general way, $\Omega_{k}^{\mathsf{g}}\neq 0$ implies the existence of a Dk1D-k-1-global symmetry in our theory, with associated conserved topological charge/invariant McNamara and Vafa (2019)
@@
where $G_{k}$ is some associated closed $k$-form. Note that by definition $Q$ is independent of the bordism representative. Now, in order to kill this bordism invariant, a Dk1D-k-1 defect is needed. For the case of oriented bordisms in the presence of $A_{p}$ forms, $\Omega_{p+1}^{SO,\,A_{p}}\simeq\mathbb{Z}$, represented by the magnetic charge $Q$ above. The D(p+1)1=Dp2D-(p+1)-1=D-p-2 defects are given by the D(Dp3)(D-p-3)-branes on whose world-volume the magnetic dual gauge potential $\tilde{A}_{D-p-2}$ lives (see McNamara and Vafa (2019) for more details on this).
@@
Figure 5: Illustrations of thin and thick walls in different bordisms, respectively for $\mathcal{C}_{2}=\mathbb{T}^{2}$ (subfigure 5(a)) and $\mathcal{C}_{1}=\mathbb{S}^{1}$ (subfigure 5(b)). In the thin wall case the (d+nk1=d)(d+n-k-1=d)-defect simply extends over d1d-1 spacetime directions and wrapping along the (nk=21=1)(n-k=2-1=1)-cycle $\tilde{a}_{1}$ dual to the 1-cycle $a_{1}$ that is shrinking. In the thick wall case the analogous $d$-defect extends over d1d-1 spacetime directions and wraps a 1-cycle $A_{1}$ of $\mathcal{B}_{2}\simeq\mathbb{T}^{2}-\mathbb{D}^{2}$. This second example can be found in section 7.3 from Debray et al. (2023).
@@
In general, when compactifying on $\mathcal{C}_{n}$ to d=Dnd=D-n dimensions, we will also have internal $k$-cycles over which topological charges can be defined. The d+nk1d+n-k-1-defects associated with $\{\Omega_{k}^{\mathsf{g}}\}_{k=0}^{n}$ will allow the associated $k$-cycles to go to nothing in the intermediate topology changes described in section 3.2. As at the different sides of the topology change (3.1) the compactification manifold $\mathcal{C}_{n}^{(k)}\to\mathcal{C}_{n}^{(k+1)}$ will be different, the lower $d$-dimensional EFT will change, with the d+nk1d+n-k-1-defects being seen from the macroscopic point of view as codimension-1 d1d-1-domain walls, which are localized at the topology change loci. Depending on how the additional nkn-k directions are wrapped, we can have the following two types of domain walls:
@@
  • Thin domain walls: The d+nk1d+n-k-1-defect extends over d1d-1 of the macroscopic directions, while wrapping the nkn-k-cycle dual to the $k$-cycle in $\mathcal{C}_{n}$ supporting the topological charge, and that the defect allows to shrink to nothing. Note that after the topology change the dual nkn-k-cycle no longer exists either, so the domain wall also acts as a boundary of both cycles. An illustration of this is depicted in Figure 5(a), and will be found in explicit examples in section 4.2.
@@
  • Thick domain walls: The d+nk1d+n-k-1-defect extends over d1d-1 macroscopic directions (perpendicular to the bordism radial direction $\rho$), while wrapping a nkn-k-cycle in the bordism $\mathcal{B}_{n+1}$, in such a way that the cycle also extends in the $\rho$ direction (for a finite interval), with the nkn-k-cycle not corresponding to any specific cycle in any of the $\mathcal{C}_{n}^{(k)}$ slices. Note that in this case the localization occurs around said cycle, which by definition is non-contractible, unlike those of $\mathcal{C}_{n}^{(k)}$ once embedded in $\mathcal{B}_{n+1}$.
@@
    This is illustrated in Figure 5(b), with an example for this disposition being found in section 7.3 of Debray et al. (2023) (albeit not in the context of an explicit BoN construction), the non-trivial 1-cycle $A_{1}\subset\mathcal{B}_{2}$ has a monodromy that kills the $\mathbb{Z}_{3}$ factor of $\Omega_{1}^{{\rm Spin-Mp}(2,\mathbb{Z})}\simeq\mathbb{Z}_{8}\oplus\mathbb{Z}_{3}$ by performing a reflection $g\to g^{-1}$ on the generator of $\mathbb{Z}_{3}=\langle g\rangle$ (we refer to the original, Debray et al. (2023) for more details). This is equivalent to having a 8-defect spanning between the two critical points (in black in Figure 5(b)) associated to $A_{1}$, resulting in a thick (7+1)-domain wall. [^13] We thank Markus Dierigl and Miguel Montero for referencing a specific example of this class to us.
@@
The defects considered here in general are of stringy origin and will backreact on the spacetime geometry, resulting in a loss of the smooth geometry (e.g. conical defects and angular deficits). We expect that, in some of these cases, the topological description of our bordism holds in a way that can be smoothed-out and the Morse-Bott inequalities hold.